OpenStudy (anonymous):

solve: lim(x->pie/2) sinx^(tanx)

OpenStudy (diyadiya):

\[\lim_{x \rightarrow \frac{\pi}{2}} sinx^{tanx}=\sin({ \frac{\pi}{2}})^\tan \frac{\pi}{2}= 1\]

OpenStudy (anonymous):

i don't understand how you got 1 ans??